# Enhancing Question Answering on Charts Through Effective Pre-training Tasks

### v1

@@ Abstract
To completely understand a document, the use of textual information is not enough. Understanding visual cues, such as layouts and charts, is also required. While the current state-of-the-art approaches for document understanding (both OCR-based and OCR-free) work well, a thorough analysis of their capabilities and limitations has not yet been performed. Therefore, in this work, we addresses the limitation of current VisualQA models when applied to charts and plots. To investigate shortcomings of the state-of-the-art models, we conduct a comprehensive behavioral analysis, using ChartQA as a case study. Our findings indicate that existing models particularly underperform in answering questions related to the chart’s structural and visual context, as well as numerical information. To address these issues, we propose three simple pre-training tasks that enforce the existing model in terms of both structural-visual knowledge, as well as its understanding of numerical questions. We evaluate our pre-trained model (called MatCha-v2) on three chart datasets - both extractive and abstractive question datasets - and observe that it achieves an average improvement of 1.7% over the baseline model.
@@ 2.1 Checklist For ChartQA
To perform the perturbation analysis for the chart- based models, we choose the popular ChartQA dataset, where the objective is to answer questions based on the information provided in the accompanying charts. This real-world dataset consists of four (4) different subsets: (a) Statista-H, (b) Pew, (c) Our World In Data (OWID), and (d) OECD. However, in all our checklist analysis and evaluation, we only use one subset of data, namely OWID, as the library (owid-grapher) to generate the charts and apply perturbations is available [^1] https://github.com/owid/owid-grapher-py. All the other data subsets were either manually curated or hard-coded and hence cannot be perturbed. To construct the checklist for behavioral analysis, we leverage 400 different charts and the corresponding data points sourced from Our World In Data. To ensure the accuracy and reliability of our analysis, we manually design three distinct types of templates: Structural & Visual, Data Extraction, and Numerical QA.
@@ 2.2 Evaluation and Results
Our focus in this work are the two large, recently introduced chart pre-trained models in MatCha, and DePlot + LLM. While MatCha is an end-to-end chart-to-text pre-trained model, DePlot + LLM is a pipelined approach where DePlot first converts an input chart into its textual representation in the form of a table and then performs few-shot question answering via a Large Language Model (LLM). In our experiments, we use the flan-ul2 (Tay et al., 2022) with 20 billion parameters as the LLM for DePlot + LLM evaluation. We do not evaluate models such as VisionTapas (Masry et al., 2022), Vl-T5 (Cho et al., 2021) since they are harder to work with than the two models we use.
@@
For each of the templates we use, we measure the model’s Failure Rate, i.e. the number of examples where model prediction does not match the expected/gold output. We present examples of some failure cases along with failure rates in Table 1.
@@ Results
We also notice intriguing disparities in the models’ performance concerning numerical QA templates. While the failure rates soar for complex mathematical operations, such as intricate calculations involving multiple operators, a marked improvement is observed in tasks involving simpler operations like finding maximum or minimum values. This nuanced discrepancy suggests that while the models struggle with advanced numerical reasoning, they exhibit a more stable grasp on elementary mathematical concepts. This insight not only sheds light on the specific challenges faced by these models in handling complex mathematical operations but also underscores their potential strengths in addressing simpler, more straightforward numerical queries.
@@ 3.1 Proposed Pre-training Tasks
We pre-train both models on the charts extracted from the training data of the ChartQA dataset. We pre-train each model for only one epoch as we saw significant reduction in accuracy on the validation set. In addition, we use batch size of six (6) and use all the other hyperparameters as suggested by  (Liu et al., 2023). We denote our proposed versions of MatCha and DePlot with the v2 suffix.
Computing Infrastructure Used
All of our experiments required access to GPU accelerators. We ran our experiments on two machines: NVIDIA Tesla V100 (16 GB VRAM) and Tesla V100 (32 GB VRAM).
@@ 3.3 Evaluation Metrics and Datasets
We evaluate the effectiveness of our proposed pre-training methods on the three chart related datasets. Here is a brief summary of each dataset along with how we use it to evaluate a model’s chart understanding capabilities:
Chart QA The Chart Question Answering (QA) (Masry et al., 2022) dataset, as the name suggests, is a natural language query and answer generation dataset with one key difference from most NLP datasets – it also provides the visual representation of data or the chart image which contain richer information such as layout, colors, etc. It contains approximately 28,000 training examples and comes with two evaluations sets: Augmented and Human. The Augmented set was constructed using a question generation system, while the Human set was made entirely from human annotations.
@@ 3.4 Results with Pre-training
As can be observed from Table  2, we see consistent improvements for the MatCha model across all three evaluation datasets. Perhaps surprisingly, we observe a much larger improvement for PlotQA than for the ChartQA dataset from where pre-training data was used. We also gain improvements for the OpenCQA dataset, which requires long-form question answering. The improvement for DePlot-v2 is smaller than that for MatCha-v2, as the component responsible for question answering is the LLM that remains unchanged.
@@ Appendix A Appendix
A.1 Implementation Details
We use batch size of six (6) to train the MatCha models due to computational constraints. Additionally, all the models that require training (e.g., MatCha) were trained up to five epochs due to computational constraints.
We did not experiment with VisionTapas (Masry et al. (2022)) as we could not run the publicly released implementation due to a missing dependency. We created a github issue that was not addressed  [^2] https://github.com/vis-nlp/ChartQA/issues/9.
We train all our models using the Transformers library from Hugging Face (Wolf et al., 2019) with the PyTorch back-end (Paszke et al., 2019).
@@ A.1 Implementation Details
[floatPos=p,capWidth=1,capPos=t,rotAngle=90,objectPos=l]table Chart Capability & Template Description Failure Rate (%) Example with Prediction (P) and Gold (G) MatCha DePlot MatCha-v2 DePlot-v2 Structural & Visual Colors in Chart: Is a certain color present or absent? 98.9 99.5 15.6 23.5 Q: Is there a bar of orange color in the given chart? P: yes, G yes Q: Is there a bar of blue color in the given chart? P: yes, G: no Chart Type: Is it bar plot or a line plot? 99.4 74.9 2.2 8.4 Q: Is it a bar plot or line plot? P: line plot, G: bar Data Extraction Extract Entity Name from Original Chart (a) 33.9 1.6 31.2 1.8 Q: Which country has wasting percentage as 9.8? P: Fiji, G: Fiji Extract Entity Value from Original Chart (a) 63.5 0.8 25.6 1.1 Q: What is the wasting percentage of Fiji in the given chart? P: 3.8, G: 9.8 Extract Value from Perturbed Chart 13.3 1.6 12.4 1.5 Q: What is the wasting percentage of Fiji in the given chart? - Sort Descending Order (b) P: 0.8, G: 9.8 Extract Value from Perturbed Chart 35.0 1.1 31.2 1.1 Q: What is the wasting percentage of Fiji in the given chart? - Add Irrelevant Bar (c) P: 0.8, G: 9.8 Numerical QA Operator: Sum 99.5 45.8 62.3 44.5 Q: What is the sum of wasting percentage for Nigeria and Sri Lanka bars? P: 23.5, G: 35.6 Operator: ArgMax 16.3 15.2 11.5 16.1 Q: Which country has the highest wasting percentage? P: Malnutrition, G: Lesotho Operator: Average + Comparison 83.4 65.2 47.5 61.0 Q: How many countries have wasting percentage more than average of all the countries? P: 2, G: 3 A partial selection of Template based tests for the ChartQA using our checklist. We report Failure Rate (in %) for each of the templates. The proposed v2 versions significantly decrease the failure rate.tab:app:failure

### v2

@@ Abstract
To completely understand a document, the use of textual information is not enough. Understanding visual cues, such as layouts and charts, is also required. While the current state-of-the-art approaches for document understanding (both OCR-based and OCR-free) work well, we have not found any other works conducting a thorough analysis of their capabilities and limitations. Therefore, in this work, we address the limitation of current VisualQA models when applied to charts and plots. To investigate shortcomings of the state-of-the-art models, we conduct a comprehensive behavioral analysis, using ChartQA as a case study. Our findings indicate that existing models underperform in answering questions related to the chart’s structural and visual context, and also numerical information. To address these issues, we propose three simple pre-training tasks that enforce the existing model in terms of structural-visual knowledge, and its understanding of numerical questions. We evaluate our pre-trained model (called MatCha-v2) on three chart datasets - both extractive and abstractive question datasets - and observe that it achieves an average improvement of 1.7% over the baseline model.
@@ 2.1 Checklist For ChartQA
To perform the perturbation analysis for the chart- based models, we choose the popular ChartQA dataset, where the objective is to answer questions based on the information provided in the accompanying charts. This real-world dataset consists of four (4) different subsets: (a) Statista-H, (b) Pew, (c) Our World In Data (OWID), and (d) OECD. However, in all our checklist analysis and evaluation, we only use one subset of data, namely OWID, as the library (owid-grapher) to generate the charts and apply perturbations is available. [^1] https://github.com/owid/owid-grapher-py All the other data subsets were either manually curated or hard-coded and hence cannot be perturbed. To construct the checklist for behavioral analysis, we leverage 400 different charts and the corresponding data points sourced from Our World In Data. To ensure the accuracy and reliability of our analysis, we manually design three distinct types of templates: Structural & Visual, Data Extraction, and Numerical QA.
@@ 2.2 Evaluation and Results
Our focus in this work entails the two large, recently introduced chart pre-trained models in MatCha, and DePlot + LLM. While MatCha is an end-to-end chart-to-text pre-trained model, DePlot + LLM is a pipelined approach where DePlot first converts an input chart into its textual representation in the form of a table and then performs few-shot question answering via a Large Language Model (LLM). In our experiments, we use the flan-ul2 (Tay et al., 2022) with 20 billion parameters as the LLM for DePlot + LLM evaluation. We do not evaluate models such as VisionTapas (Masry et al., 2022), Vl-T5 (Cho et al., 2021) since they are harder to work with than the two models we use.
@@
For each of the templates we use, we measure the model’s Failure Rate, i.e., the number of examples where model prediction does not match the expected/gold output. We present examples of some failure cases along with failure rates in Table 1.
@@ Results
We also notice intriguing disparities in the models’ performance concerning numerical QA templates. While the failure rates soar for complex mathematical operations, such as intricate calculations involving multiple operators, a marked improvement is observed in tasks involving simpler operations like finding maximum or minimum values. This nuanced discrepancy suggests that while the models struggle with advanced numerical reasoning, they exhibit a more stable grasp on elementary mathematical concepts. This observation not only sheds light on the specific challenges faced by these models in handling complex mathematical operations but also underscores their potential strengths in addressing simpler, more straightforward numerical queries.
@@ 3.1 Proposed Pre-training Tasks
Figure 4: Two of the proposed pre-training procedures: Visual Structure prediction and Summary Statistics prediction. In the example shown, for the pre-training task involving visual structure prediction, the model is asked to predict the color of the bar corresponding to each entity as well as the structure and type of the chart. As shown, for the Summary Statistics prediction, the model has to output the statistics of the data shown in the chart (ex: Maximum, Median, etc.).
@@
An example for two of these three pre-training tasks is shown in fig  4.
@@
We pre-train both models on the charts extracted from the training data of the ChartQA dataset. We continue pre-training from the initial pretrained variants and pre-train each model for only one epoch as we saw significant reduction in accuracy on the validation set. In addition, we use batch size of six (6) and use all the other hyperparameters as suggested by  (Liu et al., 2023a). We denote our proposed versions of MatCha and DePlot with the v2 suffix.
Implementation Details and Computing Infrastructure Used
We use batch size of six (6) to train the MatCha models due to computational constraints. Additionally, all the models that require training (e.g., MatCha) were trained up to five epochs. All of our experiments required access to GPU accelerators. We ran our experiments on two machines: NVIDIA Tesla V100 (16 GB VRAM) and Tesla V100 (32 GB VRAM). We did not experiment with VisionTapas (Masry et al. (2022)) as we could not run the publicly released implementation due to a missing dependency.  [^2] An issue on the github repository of the code base: https://github.com/vis-nlp/ChartQA/issues/9 We train all our models using the Transformers library from Hugging Face (Wolf et al., 2019) with the PyTorch back-end (Paszke et al., 2019).
@@ 3.3 Evaluation Metrics and Datasets
We evaluate the effectiveness of our proposed pre-training methods on the three chart-related datasets. Here is a brief summary of each dataset, along with comments on how we use it to evaluate a model’s chart understanding capabilities:
Chart QA The Chart Question Answering (QA) (Masry et al., 2022) dataset, as the name suggests, is a natural language query and answer generation dataset with one key difference from most NLP datasets – it also provides the visual representation of data or the chart image, which contain richer information such as layout, colors, etc. It contains approximately 28,000 training examples and comes with two evaluations sets: Augmented and Human. The Augmented set was constructed using a question generation system, while the Human set was made entirely from human annotations.
@@ 3.4 Results with Pre-training
As can be observed from Table  2, we see consistent improvements for the MatCha model across all three evaluation datasets. Perhaps surprisingly, we observe a much larger improvement for PlotQA than for the ChartQA dataset from which pre-training data was used. We also gain improvements for the OpenCQA dataset, which requires long-form question answering. The improvement for DePlot-v2 is smaller than that for MatCha-v2, as the component responsible for question answering is the LLM that remains unchanged.
@@
4 Related Work
Numerous studies have highlighted various robustness challenges in NLP models, including their over-sensitivity to minor perturbations (Ebrahimi et al., 2018; Wallace et al., 2019) as well as under-sensitivity to large changes (Gupta et al., 2021; Feng et al., 2018). A prominent method for identifying these issues is through behavioral analysis using specifically designed checklist examples (Ribeiro et al., 2020). In this work, we build upon this approach by applying it to multimodal chart-reasoning models, developing a checklist that aids in identifying similar robustness concerns.
Tackling these robustness challenges presents a distinct set of difficulties. Although increasing model size has resolved some of these issues (Gupta et al., 2024), this approach is not always ideal. In this study, we demonstrate that certain robustness problems can be mitigated by designing targeted pre-training tasks. For instance, we introduce a pre-training task where the model predicts the visual structure of the input chart, which improves its ability to answer questions related to colors and plot types. Similar to our work, UniChart (Masry et al., 2023) explores pre-training a multimodal model to perform both low-level tasks (e.g., extracting visual elements) as well as high-level tasks (e.g., chart summarization).
@@ 5 Conclusion
6 Limitations
We foresee one main limitation of this work. We do not conduct checklist analyses and evaluation with the latest proprietary models like GPT-4o (Achiam et al., 2023), or Claude 3.5 Sonnet. The main reason for this is the cost and budget restrictions of our project due to the large number of checklist-based evaluation examples. Additionally, we do not experiment with the latest open-source Large Multimodal Models (LMMs) such as LLaVA-1.5 (Liu et al., 2023b) or LLaVA-NeXT (Liu et al., 2024) as we found them to underperform the task-specific chart models like MatCha. [^3] Please refer to this google sheet for a comparison of the ChartQA numbers: https://docs.google.com/spreadsheets/d/1a5ImfdKATDI8T7Cwh6eH-bEsnQFzanFraFUgcS9KHWc/edit?gid=0
@@ Appendix A Appendix
A.1 Checklist Examples
We show the extended version of the table 1 in LABEL:tab:app:failure.
@@ A.1 Checklist Examples
[floatPos=p,capWidth=1,capPos=t,rotAngle=90,objectPos=l]table Chart Capability & Template Description Failure Rate (%) Example with Prediction (P) and Gold (G) MatCha DePlot MatCha-v2 DePlot-v2 Structural & Visual Colors in Chart: Is a certain color present or absent? 98.9 99.5 15.6 23.5 Q: Is there a bar of orange color in the given chart? P: yes, G yes Q: Is there a bar of blue color in the given chart? P: yes, G: no Chart Type: Is it bar plot or a line plot? 99.4 74.9 2.2 8.4 Q: Is it a bar plot or line plot? P: line plot, G: bar Data Extraction Extract Entity Name from Original Chart (a) 33.9 1.6 31.2 1.8 Q: Which country has wasting percentage as 9.8? P: Fiji, G: Fiji Extract Entity Value from Original Chart (a) 63.5 0.8 25.6 1.1 Q: What is the wasting percentage of Fiji in the given chart? P: 3.8, G: 9.8 Extract Value from Perturbed Chart 13.3 1.6 12.4 1.5 Q: What is the wasting percentage of Fiji in the given chart? - Sort Descending Order (b) P: 0.8, G: 9.8 Extract Value from Perturbed Chart 35.0 1.1 31.2 1.1 Q: What is the wasting percentage of Fiji in the given chart? - Add Irrelevant Bar (c) P: 0.8, G: 9.8 Numerical QA Operator: Sum 99.5 45.8 62.3 44.5 Q: What is the sum of wasting percentage for Nigeria and Sri Lanka bars? P: 23.5, G: 35.6 Operator: ArgMax 16.3 15.2 11.5 16.1 Q: Which country has the highest wasting percentage? P: Malnutrition, G: Lesotho Operator: Average + Comparison 83.4 65.2 47.5 61.0 Q: How many countries have wasting percentage more than average of all the countries? P: 2, G: 3 A partial selection of template-based tests for the ChartQA using our checklist. We report Failure Rate (in %) for each of the templates. The proposed v2 versions significantly decrease the failure rate.tab:app:failure